OpenStudy (jeffyblood):

Please help!!! I will fan and medal!! 1. Given 4x – 8y = 8: a. Transform the equation into slope-intercept form. b. Find the slope and y-intercept of the line. c. Find the equation, in point-slope form, of the line that is perpendicular to this line and passes through the point (1, 2). Show your work

OpenStudy (gniiij):

I can help

OpenStudy (jeffyblood):

Ok Just please make sure it's your work. Not someone elses

OpenStudy (gniiij):

A should be: 4x-8y=8 -8y=8-4x -8y=-4x+8 y=1/2x-1

OpenStudy (gniiij):

B should be: slope-1/2 y-intercept is -1

OpenStudy (jeffyblood):

Ok thank you what about c?

OpenStudy (gniiij):

i am not sure about c

OpenStudy (gniiij):

i can try though

OpenStudy (jeffyblood):

Ok please

OpenStudy (jeffyblood):

@DullJackel09 @just_one_last_goodbye @Samara8954

OpenStudy (gniiij):

ok

OpenStudy (jeffyblood):

@ILovePuppiesLol

OpenStudy (jeffyblood):

@johnweldon1993

OpenStudy (jeffyblood):

@myininaya